Operation Summary for November 17, 2010 - November 19, 2010 |
Wednesday Nov. 17 The day shift began with Operations monitoring the Tevatron (TeV) store 8277 (39.72E30) and the Recycler stash (310.6E10), with the Antiproton Source (Pbar) stacking (13.2E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and the Meson FTBF T-992 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, at 8:12 AM, store 8277 aborted when a TeV RF station (TRF2) tripped on a spark indication. RF personnel checked TRF1 and 2. Operations put the TeV into HEP shot setup and established store 8278 at 10 AM with an average initial luminosity (AIL) of 301.9E30. Operations turned off the T-992 beam for the installation of a secondary experiment (T-979). Operations turned on T-992 at 11:44 AM, while T-979 experimenters finished getting their necessary signoffs to take beam. At 2:50 PM, store 8278 quenched due to a separator spark. During the evening shift, operators completed the TeV turn-on and checkout at 9:25 AM. Operations established store 8282 at 10:30 PM. Thursday Nov. 18 The midnight shift began with Operations monitoring store 8282 (194.4E30) and the stash (64.9E10), with Pbar stacking (12.7E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations monitored the complex. During the day shift, Operations terminated store 8282 at 3:31 PM and put the TeV into shot setup. During the evening shift, Operations established store 8283 at 4:36 PM with an AIL of 336.19E30. Friday Nov. 19 The midnight shift began with Operations monitoring store 8283 (120.24E30) and the stash (214.9E10), with Pbar stacking (7E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 2:07 AM, store 8283 quenched due to a separator spark. Operations recovered the TeV cryo system, conducted a turn-on and checkout, and put the TeV into shot setup. Operations established store 8285 at 6:15 AM with an AIL of 317.9E30. The Plans for Friday and the Weekend The plans for today and the Weekend are to stack, stash, and store. The Run Coordinator will allow MI and Booster personnel to conduct studies. |
|
More Information Comments and Suggestions |
| last modified 11/19/2010 email Fermilab |
| Security, Privacy, Legal | |